OpenStudy (anonymous):

A clothing store has twice as many blue shirts as red shirts, and one-tenth as many yellow shirts as blue shirts. If the store has twenty red shirts, how many red, blue, and yellow shirts does the store have total?

OpenStudy (anonymous):

I got to 60, what do i do with the one-tenth?

OpenStudy (anonymous):

64! :D

OpenStudy (anonymous):

easiest way of solving is by writing of a system of equations relating everything. r=# of red shirts b=# of blue shirts y=# of yellow shirts b=2r y=.1b r=20 so b = 2r so b=2*10=20 y=.1*b = .1*20 = 2 total number of shirts = r + b + y = 10 + 2 + 20 = 32

OpenStudy (anonymous):

i apologize, my math was terrible there. i just made 20 into 10.... should be b = 2*20 = 40 y = .1 * 40 = 4 total = 64

OpenStudy (anonymous):

I don't think it's 32

OpenStudy (anonymous):

oh np, thxs anywayz!

OpenStudy (anonymous):

:D Thanks for the medal :D

OpenStudy (anonymous):

ur welcome! <3